OpenStudy (anonymous):

find (f of g)(x) and (g of f)(x) f(x)=x+3 g(x)=1+x^2

OpenStudy (lanhikari22):

well (f o g)(x) is the same as saying f(g(x)) so, (f o g) = (1+x^2) + 3 = x^2 + 4 and (g o f)(x) would be g(f(x)) you just put them on each other and replace x with the other function that acts as the parameter so, (g o f)(x) = 1 + (x+3)^2, i hope i helped.

OpenStudy (anonymous):

thanks! it did

OpenStudy (lanhikari22):

my pleasure.